# SpotFormer: Multi-Scale Spatio-Temporal Transformer for Facial Expression Spotting

### v1

@@ 2.3 Transformer
The Transformer was first proposed by Vaswani et al. [43] for natural language processing. In recent years, many researchers have explored the Transformer for computer vision and have shown great success in 2D image processing (e.g., ViT [44], Swin Transformer [45]) and 3D video understanding (e.g., ViViT [46], Video Swin Transformer [47], ActionFormer [33]). Such success is attributed to their powerful self-attention mechanism and ability to model long-range relations. The self-attention mechanism allows the model to focus on critical features and ignore irrelevant information, enhancing performance and robustness. The long-range relation modeling enables efficient capture of context information across both spatial and temporal dimensions, contributing to a better understanding of the overall sequence. In this paper, we explore the potential of the graph-based Transformer for facial expression spotting and introduce SpotFormer, a novel model designed to enhance the accuracy of expression spotting.
@@ 2.4 Contrastive learning
In recent years, contrastive learning has proven to be effective in various domains, including computer vision and natural language processing. The objective of contrastive learning in unsupervised learning [48] is to acquire discriminative representations by maximizing the similarity between similar instances (positive pairs generated using data augmentation) while minimizing the similarity between dissimilar instances (negative pairs). This approach produces meaningful visual representations applicable to tasks such as image recognition, representation learning, and semantic understanding. Based on the idea that positive pairs can also be selected by ground-truth labels, Khosla et al. [49] introduced the supervised contrastive loss, showcasing its potential to enhance supervised tasks by incorporating labeled information during the training process. Supervised contrastive learning efficiently enlarges the domain discrepancy, leading to an enhanced extraction of discriminative feature representations. Our method focuses on recognizing the boundary between different types of expressions. To achieve this objective, we establish contrasts between MaE and ME frames, as well as between ME frames and neutral frames. This approach enables our model to acquire more discriminative feature representations, ultimately reducing the misclassification rate.
@@ 3 Proposed Expression Spotting Framework
Figure 4: Overview of the proposed SpotFormer. (a) The network structure of SpotFormer; (b) the scale change between different facial graph structures through FLGP; (c) the illustration of the spatial graph attention layer and temporal node attention layer.
@@ 3.1 Data pre-processing
Assuming that a video $V=(v_{i})_{i=1}^{N}$ with $N$ frames, we initially calculate the SW-MRO features. Specifically, by setting a window length $w$, we pad the beginning and end of the video with $\lfloor\frac{w}{2}\rfloor$ repetitions of the first and end frames, respectively. Then, we divide the entire video into $N$ overlapped clips $C=(c_{i})_{i=1}^{N}$ with the window length $w$ and sliding stride 1, where $c_{i}$ involves $v_{i}$ as the center frame. In this paper, we aim to achieve accurate frame-level probability estimation based on compact sliding windows. Specifically, we utilize all the temporal information of $c_{i}$ and magnify the motion information existing in $c_{i}$ to estimate the probability map of $v_{i}$. Next, we will introduce the process of obtaining multi-resolution optical flow features in each compact sliding window.
For each clip $c_{i}$, we initially detect the 68 facial keypoints using a pre-trained MobileFaceNet [50] in the first frame $c_{i}^{1}$, which are then utilized for face cropping, extracting the nose tip area for face alignment, and extracting ROIs. Specifically, we detect the facial bounding box in the first frame $c_{i}^{1}$. In each subsequent frame $c_{i}^{s}$, where $s=2,3,\ldots,w$, we initialize the facial bounding box with the one from $c_{i}^{1}$. Given that the nose tip area remains stationary during expressions [13], we compute the optical flow of the nose tip area, as illustrated in Fig. 3, to represent global head movement. Subsequently, we adjust the facial bounding box for $c_{i}^{s}$ accordingly. To achieve this, we employ the Farneback algorithm [51] to compute MDMO optical flow [32] of the nose tip area $o_{i}^{s,\mathrm{nose}}\in\mathbb{R}^{2}$ between $c_{i}^{1}$ and $c_{i}^{s}$:
@@
where $m^{\mathrm{nose}}_{i}$ and $n^{\mathrm{nose}}_{i}$ represent the height and width of the extracted nose tip area $M^{\mathrm{nose}}_{i}$, and $\mathrm{OF}(\cdot,\cdot)[x,y]$ denotes the value of MDMO optical flow features at $(x,y)$ coordinates.
Based on the discussion in Sections 1 and 2, we selectively extract $R$ ROIs, where facial expressions happen most frequently, rather than processing entire images. This choice can effectively mitigate the impact of noise and irrelevant facial movements. Then, we compute the optical flow for the selected $R$ ROIs between $c_{i}^{1}$ and $c_{i}^{s}$ to obtain MDMO optical flow features $o_{i}^{s}=[o_{i}^{s,r}]_{r=1}^{R}\in\mathbb{R}^{R\times 2}$. The computation is similar to (1). Afterwards, we concatenate the optical flow $[o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{w}]$ to construct finer optical flow features $o_{i}\in\mathbb{R}^{w\times R\times 2}$ for $c_{i}$, where $o_{i}^{1}=\bm{0}$. As a result of the data pre-processing, we obtain the optical flow features $O=(o_{i})_{i=1}^{N}\in\mathbb{R}^{N\times w\times R\times 2}$ for the entire input video.
@@ 3.2 SpotFormer for probability estimation
The overview of the proposed SpotFormer is outlined in Fig. 4. After obtaining the optical flow features, the SpotFormer calculates the probabilities required for expression spotting. These include the probabilities of occurrence for onset, apex, offset, expression, and neutral states in both ME and MaE spotting. SpotFormer consists of graph node embedding, temporal positional embedding, spatial facial embedding, and multi-scale spatio-temporal attention block, aiming to comprehensively learn both spatial relations and temporal variations for accurate estimation of the probability map of every single frame. Subsequently, we will introduce each component in detail.
@@ 3.2.2 Temporal positional embedding
For the temporal dimension, following [44], we set tokens with learnable parameters, denoted as $\delta_{\mathrm{t}}\in\mathbb{R}^{w\times D}$, to retain temporal positional information. Note that all graph patches in a single frame share the same temporal positional embedding.
@@ 3.2.3 Spatial facial embedding
Unlike other graph-based tasks where there are predefined edge connections among graph nodes, facial expression analysis involves many possible edge connections among different facial muscles. To mitigate the impact of these various sophisticated edge connections, SpotFormer does not explicitly consider the edges but focuses on the facial part type of each graph node instead. Inspired by [52], we introduce learnable facial part tokens to the graph nodes. Specifically, we set three tokens with learnable parameters $\delta_{\mathrm{s}}\in\mathbb{R}^{3\times D}$ for three asymmetric facial parts (left eyebrow, right eyebrow, and mouth) since MEs are asymmetric. Graph nodes belonging to the same facial part share the same facial part token. This method enables the model to understand the overall spatial facial structure without the requirement of exploring complex edge connections.
@@ 3.2.5 Multi-scale spatio-temporal feature learning
Multi-scale learning has shown powerful performance in image processing [53, 45] and video understanding [54, 42]. It is also significant in facial expression analysis since it enables the model to extract both coarse and fine feature representations across different scales, enhancing its ability to capture comprehensive facial structures and temporal dynamics.
For extracting multi-scale temporal features, we use a 1D CNN with stride $2$ to achieve temporal downsampling, following ActionFormer [33]. However, for spatial downsampling, applying pooling operations, which are generally used for downsampling images, to graph-structured data presents challenges because it is a type of non-Euclidean structured data. To address this issue, inspired by Xu et al. [55], we introduce FLGP, specifically designed for extracting multi-scale facial graph features. In practice, we design three scales of facial structures. The designed scales and the scale change achieved through FLGP are illustrated in Fig. 4 (b). During each FLGP operation, the facial graph is downsampled by aggregating features from multiple nodes using the max pooling operation.
@@ 3.2.6 Model architecture
SpotFormer simultaneously models spatio-temporal relations, as shown in Fig. 4 (a). Specifically, in each spatio-temporal attention block, the model first employs a spatial graph attention layer to model intricate spatial relationships, followed by a temporal node attention layer that captures temporal variations among each graph node across frames. After each spatio-temporal attention block, spatial downsampling and temporal downsampling are performed to aggregate information to generate high-level features. Note that when the facial graph structure is downsampled to one node, subsequent temporal information learning relies solely on temporal node attention layers and temporal downsampling until the temporal resolution is reduced to 1.
@@ 3.2.7 Frame-level probability estimation
Finally, after reducing the spatial graph to a single node and temporal information to a single frame through SpotFormer, the remaining single node comprehensively aggregates spatio-temporal information within a compact sliding window. We then use a fully-connected (FC) layer to generate the probability map $p_{i}=\{p_{i}^{\mathrm{onset}}$, $p_{i}^{\mathrm{apex}}$, $p_{i}^{\mathrm{offset}}$, $p_{i}^{\mathrm{exp}}$, $p_{i}^{\mathrm{norm}}$} corresponding to the frame $v_{i}$. This map contains the probabilities of $v_{i}$ being an onset frame, apex frame, offset frame, expression frame, or neutral frame. Additionally, each component in $p_{i}$ incorporates two probabilities for ME spotting and MaE spotting, respectively. Specifically, $p_{i}^{\mathrm{onset}}=\{p_{i}^{\mathrm{mi},\mathrm{onset}},p_{i}^{\mathrm{ma},\mathrm{onset}}\}$, $p_{i}^{\mathrm{apex}}=\{p_{i}^{\mathrm{mi},\mathrm{apex}},p_{i}^{\mathrm{ma},\mathrm{apex}}\}$, $p_{i}^{\mathrm{offset}}=\{p_{i}^{\mathrm{mi},\mathrm{offset}},p_{i}^{\mathrm{ma},\mathrm{offset}}\}$, $p_{i}^{\mathrm{exp}}=\{p_{i}^{\mathrm{mi},\mathrm{exp}},p_{i}^{\mathrm{ma},\mathrm{exp}}\}$, $p_{i}^{\mathrm{norm}}=\{p_{i}^{\mathrm{mi},\mathrm{norm}},p_{i}^{\mathrm{ma},\mathrm{norm}}\}$. The optimization tasks are divided into two binary classification tasks and a three-class classification task for different types of frames, following the optimization method outlined in [23]. Focal-loss [56] is employed to optimize our model, which can be expressed as:
@@ 3.3 Supervised contrastive learning
We have been focusing on minimizing the divergence between predicted class probabilities and ground-truth class labels so far, potentially overlooking distributional differences among various classes. We notice that distinguishing certain MaEs and MEs near the boundary poses challenges in terms of duration and intensity. Specifically, the annotation of expressions follows the criterion that MaEs are longer than 0.5 seconds, while MEs are below 0.5 seconds. This criterion introduces difficulties in distinguishing expressions whose duration is close to 0.5 seconds. Similar challenges occur when distinguishing between ME frames and neutral frames. This is due to the fact that certain ground-truth MEs exhibit very low intensity, as well as the presence of noise in the optical flow features, causing the misclassification of specific neutral frames as ME frames.
To overcome this issue, we introduce supervised contrastive learning [49] to enhance our model’s discriminative feature learning, aiming to improve its ability to recognize the boundary between different types of expressions. Specifically, we utilize the feature representation before the last classification FC layer as the feature representation for each frame. The supervised contrastive loss is then employed to minimize the distance between feature representations of the same expression class while simultaneously pushing apart feature representations of different expression classes. The expression type label $\widetilde{y}_{i}$ of the frame $v_{i}$ for the $i$-th sliding window is used as the supervision information for the supervised contrastive loss. This means that each frame is labeled as a MaE frame, ME frame, or neutral frame. Let $I$ denote a set of samples in a batch, and the loss function can be expressed as:
@@
Traditional methods MDMD[12] 0.0629 0.0364 0.0445 0.1196 0.0082 0.0376
Optical Strain[11] - - - 0.1436 0.0098 0.0448
Zhang et al.[57] 0.0725 0.1331 0.0999 0.2131 0.0547 0.1403
He [13] 0.4149 0.2162 0.3638 0.3782 0.1965 0.3436
Zhao et al.[14] - - 0.3863 - - 0.4030
Wang et al.[58] 0.3724 0.2866 0.3419 0.5061 0.2614 0.4558
Deep-learning methods Verburg[15] - 0.0821 - - - -
LBCNN[59] - - 0.0813 - - 0.0595
MESNet[16] - 0.0880 - - 0.0360 -
SOFTNet[30] 0.2169 0.1520 0.1881 0.2410 0.1173 0.2022
3D-CNN[17] 0.1595 0.0466 0.1084 0.2145 0.0714 0.1675
Concat-CNN[31] 0.3553 0.1155 0.2736 0.2505 0.0153 0.2019
LSSNet[22] 0.2810 0.1310 0.2380 0.3770 0.0420 0.3250
MTSN[60] 0.3459 0.0878 0.2867 0.4104 0.0808 0.3620
ABPN[23] 0.3349 0.1689 0.2908 0.3357 0.1590 0.3117
AUW-GCN[18] 0.4293 0.1984 0.3728 0.4235 0.1538 0.3834
LGSNet[19] - - 0.3880 - - 0.4360
MULT[20] - 0.2770 - - 0.1373 -

### v2

@@ 2.3 Transformer
The Transformer was first proposed by Vaswani et al. [transformer] for natural language processing. In recent years, many researchers have explored the Transformer for computer vision and have shown great success in 2D image processing (e.g., ViT [dosovitskiy2021an], Swin Transformer [liu2021swin]) and 3D video understanding (e.g., ViViT [Arnab_2021_ICCV], Video Swin Transformer [Liu_2022_CVPR], ActionFormer [zhang2022actionformer]). Such success is attributed to their powerful self-attention mechanism and ability to model long-range relations. The self-attention mechanism allows the model to focus on critical features and ignore irrelevant information, enhancing performance and robustness. The long-range relation modeling enables efficient capture of context information across both spatial and temporal dimensions, contributing to a better understanding of the overall sequence. In this paper, we explore the potential of the graph-based Transformer for facial expression spotting and introduce SpotFormer, a novel model designed to enhance the accuracy of expression spotting.
@@ 2.4 Contrastive learning
In recent years, contrastive learning has proven to be effective in various domains, including computer vision and natural language processing. The objective of contrastive learning in unsupervised learning [chen2020simple] is to acquire discriminative representations by maximizing the similarity between similar instances (positive pairs generated using data augmentation) while minimizing the similarity between dissimilar instances (negative pairs). This approach produces meaningful visual representations applicable to tasks such as image recognition, representation learning, and semantic understanding. Based on the idea that positive pairs can also be selected by ground-truth labels, Khosla et al. [khosla2020supervised] introduced the supervised contrastive loss, showcasing its potential to enhance supervised tasks by incorporating labeled information during the training process. Supervised contrastive learning efficiently enlarges the domain discrepancy, leading to an enhanced extraction of discriminative feature representations. Our method focuses on recognizing the boundary between different types of expressions at the frame level. To achieve this objective, we establish contrasts between MaE and ME frames, as well as between ME frames and neutral frames. This approach enables our model to acquire more discriminative feature representations, ultimately reducing the misclassification rate.
@@ 3 Proposed Expression Spotting Framework
Figure 4: Overview of the proposed SpotFormer. (a) The network structure of SpotFormer. The dotted gray box provides an intuitive illustration of the spatio-temporal attention block. A spatial graph attention layer models the spatial relationships within the graph of each frame, while a temporal node attention layer captures temporal variations of each graph node across frames. Subsequently, spatial and temporal downsampling are performed using FLGP and a learnable temporal downsampling layer, respectively. (b) The scale change between different facial graph structures through FLGP. Each facial muscle is represented by a distinct color (e.g., orange represents the lips). In the first spatial downsampling, ROIs belonging to the same facial muscle are aggregated into single nodes via FLGP to construct graph #2. In the second spatial downsampling, all aggregated nodes are further merged into a single node via FLGP to form graph #3.
(c) The illustration of the spatial graph attention layer and temporal node attention layer.
@@ 3.1 Data pre-processing
Assuming that a video $V=(v_{i})_{i=1}^{N}$ with $N$ frames, we initially calculate the SW-MRO features. First, by setting a window length $w$, we pad the beginning and end of the video with $\lfloor\frac{w}{2}\rfloor$ repetitions of the first and last frames, respectively. This ensures that the original first and last frames can be the center frames of the first and last sliding windows, respectively. Then, we divide the entire video into $N$ overlapped clips $C=(c_{i})_{i=1}^{N}$ with the window length $w$ and sliding stride 1, where $c_{i}$ involves $v_{i}$ as the center frame. In this paper, we aim to achieve accurate frame-level probability estimation based on compact sliding windows. Specifically, we utilize all the temporal information of $c_{i}$ and magnify the motion information existing in $c_{i}$ to estimate the probability map of $v_{i}$. Next, we will introduce the process of obtaining multi-temporal-resolution optical flow features in each compact sliding window.
For each clip $c_{i}$, we initially detect the 68 facial keypoints using a pre-trained MobileFaceNet [PFL] in the first frame $c_{i}^{1}$. These keypoints will subsequently be used to generate the facial bounding box, extract the nose tip area for facial alignment, and extract ROIs. Specifically, we detect the facial bounding box in the first frame $c_{i}^{1}$. In each subsequent frame $c_{i}^{s}$, where $s=2,3,\ldots,w$, we initialize the facial bounding box of $c_{i}^{s}$ with the one from $c_{i}^{1}$. Given that the nose tip area remains stationary during facial expressions [yuhong2021research], we compute the optical flow of the nose tip area, as illustrated in Fig. 3, to represent the global head movement from $c_{i}^{1}$ to $c_{i}^{s}$. To achieve this, we employ the Farneback algorithm [farneback2003two] to compute MDMO optical flow [liu2015main] of the nose tip area $o_{i}^{s,\mathrm{nose}}\in\mathbb{R}^{2}$ from $c_{i}^{1}$ to $c_{i}^{s}$:
@@
where $m^{\mathrm{nose}}_{i}$ and $n^{\mathrm{nose}}_{i}$ represent the height and width of the extracted nose tip area $M^{\mathrm{nose}}_{i}$, and $\mathrm{OF}(\cdot,\cdot)[x,y]$ denotes the value of MDMO optical flow features at $(x,y)$ coordinates. Consequently, $o^{s,\mathrm{nose}}_{i}$ contains the horizontal and vertical movements of the head from $c_{i}^{1}$ to $c_{i}^{s}$. Then, we translate the facial bounding box of frame $c_{i}^{s}$ by adding $o^{s,\mathrm{nose}}_{i}$ to complete facial alignment.
Based on the discussion in Sections 1 and 2, we selectively extract $R$ ROIs [zhao2022rethinking], where facial expressions happen most frequently, rather than processing entire images. This choice can effectively mitigate the impact of noise and irrelevant facial movements. Then, we compute the optical flow for the selected $R$ ROIs between the first frame $c_{i}^{1}$ and frame $c_{i}^{s}$ to obtain MDMO optical flow features $o_{i}^{s}=[o_{i}^{s,r}]_{r=1}^{R}\in\mathbb{R}^{R\times 2}$. The computation is similar to Eq. (1). Afterwards, we concatenate the optical flow $o_{i}^{1},o_{i}^{2},\ldots,o_{i}^{w}]$ to construct multi-temporal-resolution optical flow features $o_{i}\in\mathbb{R}^{w\times R\times 2}$ for the clip $c_{i}$, where $o_{i}^{1}=\bm{0}$. As a result of the data pre-processing, we concatenate the multi-temporal-resolution optical flow from all clips and obtain the sliding window-based multi-temporal-resolution optical flow (SW-MRO) features $O=(o_{i})_{i=1}^{N}\in\mathbb{R}^{N\times w\times R\times 2}$ for the entire input video.
@@ 3.2 SpotFormer for probability estimation
The overview of the proposed SpotFormer is outlined in Fig. 4. After obtaining the optical flow features, SpotFormer calculates the frame-level probabilities required for expression spotting. These include the probabilities of occurrence for onset, apex, offset, expression, and neutral states for each frame in both ME and MaE spotting. SpotFormer consists of graph node embedding, temporal positional embedding, spatial facial embedding, and multi-scale spatio-temporal attention block. All these modules enable SpotFormer to comprehensively learn both spatial relations and temporal variations, facilitating accurate estimation of the probability map for every single frame. Subsequently, we will introduce each component in detail.
@@ 3.2.2 Temporal positional embedding
For the temporal dimension, following [dosovitskiy2021an], we set tokens with learnable parameters, denoted as $\delta_{\mathrm{t}}\in\mathbb{R}^{w\times D}$, to retain temporal positional information. Note that all graph patches in a single frame share the same temporal positional embedding.
@@ 3.2.3 Spatial facial embedding
Unlike other graph-based tasks where there are predefined edge connections among graph nodes, facial expression analysis involves many possible edge connections among different facial muscles. To mitigate the impact of complex and diverse edge connections, SpotFormer constructs parameterized associations between nodes based on facial muscle characteristics, unlike traditional GCNs that rely on pre-defined edge connections. Inspired by [zhang2021STST], we introduce learnable facial part tokens to the graph nodes. Specifically, we set three tokens with learnable parameters $\delta_{\mathrm{s}}\in\mathbb{R}^{3\times D}$ for three asymmetric facial parts (left eyebrow, right eyebrow, and mouth) since MEs are asymmetric. Graph nodes belonging to the same facial part share the same facial part token. This method enables the model to understand the overall spatial facial structure without the requirement of exploring complex pre-defined edge connections.
@@ 3.2.5 Multi-scale spatio-temporal feature learning
Multi-scale learning has shown powerful performance in image processing [lin2017feature, liu2021swin] and video understanding [Fan_2021_ICCV, Shao_2023_ICCV]. It is also significant in facial expression analysis since it enables the model to extract both coarse and fine feature representations across different scales, enhancing its ability to capture comprehensive facial structures and temporal dynamics.
SpotFormer extracts multi-scale spatio-temporal features, enabling more comprehensive learning of motion patterns from low-level to high-level. Specifically, to extract multi-scale temporal features, we use 1D CNNs with a stride of 2 for temporal downsampling, as proposed in ActionFormer [zhang2022actionformer].
However, for spatial downsampling, applying pooling operations, which are generally used for downsampling images, to graph-structured data presents challenges because it is a type of non-Euclidean structured data. To address this issue, inspired by Xu et al. [xu2021graph], we introduce FLGP, specifically designed for extracting multi-scale facial graph features. In practice, we design three scales of facial structures. The designed scales and the scale change achieved through FLGP are illustrated in Fig. 4 (b). During each FLGP operation, the facial graph is downsampled by aggregating features from multiple nodes using the max pooling operation.
@@ 3.2.6 Model architecture
SpotFormer simultaneously models spatio-temporal relations, as shown in Fig. 4 (a). Specifically, in each spatio-temporal attention block, the model first employs a spatial graph attention layer to model intricate spatial relationships, followed by a temporal node attention layer that captures temporal variations among each graph node across frames. After each spatio-temporal attention block, spatial downsampling and temporal downsampling are performed to aggregate information to generate high-level features. Note that when the facial graph structure is downsampled to one node, subsequent temporal information learning relies solely on temporal node attention layers and temporal downsampling until the temporal dimension is reduced to 1.
@@ 3.2.7 Frame-level probability estimation
Finally, after reducing the spatial graph to a single node and temporal dimension to 1 through SpotFormer, the remaining single node comprehensively aggregates spatio-temporal information within a compact sliding window. We then use a fully-connected (FC) layer to generate the probability map $p_{i}=\{p_{i}^{\mathrm{onset}}$, $p_{i}^{\mathrm{apex}}$, $p_{i}^{\mathrm{offset}}$, $p_{i}^{\mathrm{exp}}$, $p_{i}^{\mathrm{neut}}$} corresponding to the frame $v_{i}$. This map contains the probabilities of $v_{i}$ being an onset frame, apex frame, offset frame, expression frame, or neutral frame. Additionally, each component in $p_{i}$ incorporates two probabilities for ME spotting and MaE spotting, respectively. Specifically, $p_{i}^{\mathrm{onset}}=\{p_{i}^{\mathrm{me},\mathrm{onset}},p_{i}^{\mathrm{mae},\mathrm{onset}}\}$, $p_{i}^{\mathrm{apex}}=\{p_{i}^{\mathrm{me},\mathrm{apex}},p_{i}^{\mathrm{mae},\mathrm{apex}}\}$, $p_{i}^{\mathrm{offset}}=\{p_{i}^{\mathrm{me},\mathrm{offset}},p_{i}^{\mathrm{mae},\mathrm{offset}}\}$, $p_{i}^{\mathrm{exp}}=\{p_{i}^{\mathrm{me},\mathrm{exp}},p_{i}^{\mathrm{mae},\mathrm{exp}}\}$, $p_{i}^{\mathrm{neut}}=\{p_{i}^{\mathrm{me},\mathrm{neut}},p_{i}^{\mathrm{mae},\mathrm{neut}}\}$. Therefore, the model outputs ten probabilities for each frame. The optimization tasks are divided into two groups for MaE spotting and ME spotting, respectively. Each group contains two binary classification tasks and a three-class classification task for different types of frames, following the optimization method outlined in [leng2022abpn]. Specifically, the first binary classification task determines whether the current frame is an apex frame, and the second determines whether it is an expression frame. The three-class classification task is used to classify each frame as an onset frame, an offset frame, or neither. Focal-loss [lin2017focal] is employed to optimize our model, which can be expressed as:
@@ 3.3 Supervised contrastive learning
We have been focusing on minimizing the divergence between predicted class probabilities and ground-truth class labels so far, potentially overlooking distributional differences among various classes. We notice that distinguishing certain MaEs and MEs near the boundary poses challenges in terms of duration and intensity. Specifically, the annotation of expressions follows the criterion that MaEs are longer than 0.5 seconds, while MEs are below 0.5 seconds. This criterion introduces difficulties in distinguishing expressions whose duration is close to 0.5 seconds. Similar challenges occur when distinguishing between ME frames and neutral frames. This is due to certain ground-truth MEs exhibiting very low intensity, which can be easily confused with noise in the optical flow features, leading to the misclassification of some neutral frames as ME frames.
To overcome this issue, we introduce supervised contrastive learning [khosla2020supervised] to enhance our model’s discriminative feature learning, aiming to improve its ability to recognize the classification boundaries between different types of expressions at the frame level. Specifically, we utilize the feature representation before the last classification FC layer as the feature representation for each frame. The supervised contrastive loss is then employed to minimize the distance between feature representations of the same expression class while simultaneously pushing apart feature representations of different expression classes. The expression type label $\widetilde{y}_{i}$ of the frame $v_{i}$ for the $i$-th sliding window $c_{i}$ is used as the supervision information for the supervised contrastive loss. This means that each frame is labeled as a MaE frame, ME frame, or neutral frame. Let $I$ denote a set of samples in a batch, and the loss function can be expressed as:
@@
Traditional methods MDMD[he2020spotting] 0.0629 0.0364 0.0445 0.1196 0.0082 0.0376
Optical Strain[gan2020optical] - - - 0.1436 0.0098 0.0448
Zhang et al.[zhang2020spatio] 0.0725 0.1331 0.0999 0.2131 0.0547 0.1403
He [yuhong2021research] 0.4149 0.2162 0.3638 0.3782 0.1965 0.3436
Zhao et al.[zhao2022rethinking] - - 0.3863 - - 0.4030
Wang et al.[wangunique] 0.3724 0.2866 0.3419 0.5061 0.2614 0.4558
Deep-learning methods Verburg[verburg2019micro] - 0.0821 - - - -
LBCNN[pan2020local] - - 0.0813 - - 0.0595
MESNet[wang2021mesnet] - 0.0880 - - 0.0360 -
SOFTNet[liong2021shallow] 0.2169 0.1520 0.1881 0.2410 0.1173 0.2022
3D-CNN[yap20223d] 0.1595 0.0466 0.1084 0.2145 0.0714 0.1675
Concat-CNN[yang2021facial] 0.3553 0.1155 0.2736 0.2505 0.0153 0.2019
LSSNet[yu2021lssnet] 0.2810 0.1310 0.2380 0.3770 0.0420 0.3250
MTSN[liong2022mtsn] 0.3459 0.0878 0.2867 0.4104 0.0808 0.3620
ABPN[leng2022abpn] 0.3349 0.1689 0.2908 0.3357 0.1590 0.3117
AUW-GCN[yin2023aware] 0.4293 0.1984 0.3728 0.4235 0.1538 0.3834
LGSNet[yu2023lgsnet] - - 0.3880 - - 0.4360
MULT[GUO2023146] - 0.2770 - - 0.1373 -